OpenStudy (anonymous):

What is eccentricity of the ellipse given by the equation x^2/49 + y^2/81 =1? .63 or .81

OpenStudy (amistre64):

isnt e = c/a?

OpenStudy (anonymous):

yes it is but im still confused...

OpenStudy (amistre64):

well, the major axis is the larger value; and c is the focal length

OpenStudy (anonymous):

sqrt(1-49/81)

OpenStudy (amistre64):

 they tend to be situated like this

OpenStudy (amistre64):

the equation gives us a and b; we would need to find c

OpenStudy (anonymous):

You don't need c if you have a and b

OpenStudy (anonymous):

sqrt(1-49/81) = 0.6285

OpenStudy (amistre64):

doesnt c = sqrt(a^2-b^2) ?

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Yes, it is same thing just rearranged....

OpenStudy (anonymous):

e= sqrt(1-b^2/a^2) a>b

OpenStudy (amistre64):

then finding c is just a part of the overall process. you still need to find "c" :) sqrt(a^2-b^2)/sqrt(a^2) .... simplifies to your results. yes